### v3

# Towards Clinical Practice in CT-Based Pulmonary Disease Screening: An Efficient and Reliable Framework

Abstract
Deep learning models for pulmonary disease screening from Computed Tomography (CT) scans promise to alleviate the immense workload on radiologists. Still, their high computational cost, stemming from processing entire 3D volumes, remains a major barrier to widespread clinical adoption. Current sub-sampling techniques often compromise diagnostic integrity by introducing artifacts or discarding critical information. To overcome these limitations, we propose an Efficient and Reliable Framework (ERF) that fundamentally improves the practicality of automated CT analysis. Our framework introduces two core innovations: (1) A Cluster-based Sub-Sampling (CSS) method that efficiently selects a compact yet comprehensive subset of CT slices by optimizing for both representativeness and diversity. By integrating an efficient $k$-Nearest Neighbor ($k$-NN) search with an iterative refinement process, CSS bypasses the computational bottlenecks of previous methods while preserving vital diagnostic features. (2) A lightweight Hybrid Uncertainty Quantification (HUQ) mechanism, which uniquely assesses both Aleatoric Uncertainty (AU) and Epistemic Uncertainty (EU) with minimal computational overhead. By maximizing the discrepancy between auxiliary classifiers, HUQ provides a robust reliability score, which is crucial for building trust in automated systems operating on partial data. Validated on two public datasets with $2,654$ CT volumes across diagnostic tasks for $3$ pulmonary diseases, our proposed ERF achieves diagnostic performance comparable to the full-volume analysis (over $90\%$ accuracy and recall) while reducing processing time by more than $60\%$. This work represents a significant step towards deploying fast, accurate, and trustworthy AI-powered screening tools in time-sensitive clinical settings.
Introduction
Deep learning excels at automating pulmonary disease screening using CT scans, offering a way to reduce the significant workload of radiologists (Fink et al. 2025; Hao et al. 2025; Tawfeek et al. 2025). However, the standard approach of processing entire 3D volumes is computationally expensive, hindering adoption in time-sensitive clinical settings and resource-limited environments (Balasamy and Seethalakshmi 2025; Pham et al. 2025; Zheng et al. 2024). This method expends considerable resources on diagnostically irrelevant slices, rather than quickly skimming them like radiologists do to focus on disease-relevant ones. Therefore, developing methods that intelligently select salient diagnostic information, rather than processing CT data indiscriminately, is crucial for practical clinical integration.
Figure 1: Visualization of interpolation methods on common pneumonia (CP) lesions. (a-d) are four consecutive slices showing CP lesions (red solid-line area). (e) and (f) are results of spline interpolation and projection interpolation, respectively. Yellow solid lines indicate processing-induced artifacts, blue dashed lines highlight regions of lesion elimination post-processing.
Prior work has explored sub-sampling to improve efficiency, but these methods often compromise diagnostic integrity. For instance, interpolation techniques can introduce artifacts or lose critical details (Han et al. 2023; Liauchuk 2019), as shown in Figure 1. While conventional slice selection methods may discard vital slices due to heuristic or uniform criteria (Zunair, Rahman, and Mohammed 2019; Zunair et al. 2020). Recent machine learning has introduced advanced sub-sampling strategies focusing on representativeness or diversity to tackle the above challenges (Sener and Savarese 2018; Wu et al. 2024). Representative sampling aims to select a subset whose data distribution closely mirrors that of the entire dataset, while diversity sampling endeavors to select a varied subset encompassing a broad spectrum of informative samples. Although these methods have achieved significant success in the field of machine learning, they suffer from three limitations that render them unsuitable for our task: (1) Most approaches focus solely on a single aspect. An exclusive emphasis on representativeness risks introducing redundancy and inefficiency, while prioritizing diversity alone may result in the loss of collectively important diagnostic information (Zhang et al. 2023b; Shao et al. 2024); (2) They typically require computing pairwise similarities between samples to estimate representativeness and diversity, creating a substantial computational bottleneck that contradicts our goal of efficient screening (Xie et al. 2023; Wang, Lian, and Yu 2022); (3) They fail to account for lung anatomical structure, causing significant variations in semantic information among different slices in the sub-sampled results, thereby compromising downstream task performance (Zunair et al. 2020).
This inherent challenge of ensuring diagnostic integrity within a computationally efficient sub-sampling framework elevates a second, equally critical requirement for clinical translation: robust uncertainty quantification. When a model must operate on a reduced and potentially incomplete data subset, assessing the reliability of its predictions is no longer just beneficial, but imperative for safe and trustworthy deployment. Although numerous uncertainty quantification (UQ) techniques have been introduced, including Bayesian Neural Networks (Ngartera, Issaka, and Nadarajah 2024; Vranken et al. 2021) and deep ensembles (Lakshminarayanan, Pritzel, and Blundell 2017; Smith and Gal 2018), these approaches are not readily adaptable to our problem setting. A key limitation is their typical focus on evaluating a single dimension of uncertainty: either aleatoric uncertainty (AU), originating from inherent data ambiguity (Hüllermeier and Waegeman 2021; Wang et al. 2019), or epistemic uncertainty (EU), resulting from model deficiencies (Swiler, Paez, and Mayes 2009; Nannapaneni and Mahadevan 2016). In critical applications like medical diagnostics, however, a comprehensive assessment of both AU and EU is necessary. Furthermore, the significant computational demands of these techniques make their deployment within efficient, large-scale screening pipelines infeasible (Abboud, Lombaert, and Kadoury 2024).
To address the above issues, we propose an efficient and reliable framework (ERF). For the first issue, we design a cluster-based sub-sampling (CSS) method to select representative and diverse CT slices. CSS first partitions the lung volume based on anatomical structure. Within each part, CSS uses $K$-Means to group slices (Wang, Lian, and Yu 2022) and an efficient approximate $k$-NN search algorithm based on hierarchical navigable small world graphs (HNSW) (Malkov 2018) to identify the most representative slice in each cluster. An iterative optimization process with a regularization term then refines the selection to enhance diversity, ensuring the final subset captures global characteristics while minimizing redundancy. The selected slices are used to train a lightweight model (He et al. 2016) for initial disease classification. For the second issue, we design a hybrid uncertainty quantification (HUQ) mechanism to provide a robust reliability estimate with minimal overhead. Inspired by the idea of maximizing the classifier discrepancy (Saito et al. 2018; Cho et al. 2022) in domain adaptation, we employ two auxiliary classifiers alongside the primary one. By training these auxiliary heads to maximize their predictive discrepancy on in-distribution data (Saito et al. 2018; Cho et al. 2022), we encourage them to explore diverse decision boundaries around the data manifold. This discrepancy serves as a proxy for AU. Concurrently, the prediction entropy from the primary classifier estimates EU. By integrating both UQ dimensions in a shared-backbone architecture, HUQ delivers a comprehensive uncertainty score without significant computational cost.
We validate ERF on two public datasets totaling $2,654$ CT volumes for three pulmonary diseases. Experiments demonstrate that CSS outperforms existing sub-sampling techniques under different sub-sampling rates. Furthermore, HUQ mechanism proves superior in identifying uncertain cases that correlate with misdiagnoses. Critically, ERF achieves diagnostic performance (more than $90\%$ accuracy and recall) comparable to processing full CT volumes while reducing the time by more than $60\%$. Finally, the visualization results demonstrate that CSS can select representative CT slices containing crucial indicators to aid in diagnosis.
In summary, the main contributions are as follows:
  1. 1.
    We develop ERF for pulmonary disease screening based on CT data efficiently and accurately.
  2. 2.
    We propose CSS, which selects a representative and diverse set of CT slices, significantly boosting efficiency while preserving diagnostic integrity.
  3. 3.
    We design the HUQ mechanism to detect uncertain cases, integrating both AU and EU quantification with limited additional computational overhead.
  4. 4.
    Extensive validation on public datasets demonstrating that ERF matches the accuracy and recall of full-volume analysis but reduces processing time by over $60\%$, confirming its suitability for clinical deployment.
Related Works
Pulmonary Disease Screening
The rapid development of deep learning has catalyzed a paradigm shift towards automated screening of pulmonary diseases, substantially alleviating the heavy workload of radiologists (Fink et al. 2025; Tawfeek et al. 2025). CT scans, by their superior capability in furnishing high-resolution, cross-sectional anatomical details and enhanced sensitivity in detecting subtle abnormalities compared to conventional radiography, have emerged as the imaging modality of choice for screening a wide spectrum of pulmonary diseases (Balasamy and Seethalakshmi 2025; Dally and Rekha 2024). Illustrative examples include interstitial lung diseases such as idiopathic pulmonary fibrosis (Agarwala et al. 2020; Guiot et al. 2025; Chantzi, Kosvyra, and Chouvarda 2025) and pulmonary nodules (Fink et al. 2025; Harsono, Liawatimena, and Cenggoro 2022; Jin et al. 2024); obstructive airway diseases like chronic obstructive pulmonary disease (Schroeder et al. 2020; Zhu 2024; Dally and Rekha 2024); neoplastic conditions including lung cancer (Quanyang et al. 2024; Ardila et al. 2019; Li et al. 2020); and infectious diseases such as tuberculosis (Lewis et al. 2021; Zunair et al. 2020; Zunair, Rahman, and Mohammed 2019) and pneumonia (Pham et al. 2025; Chen et al. 2025; Kordnoori et al. 2025). While most studies enhance accuracy and recall, they often overlook the bottleneck of voluminous 3D CT data, hindering widespread clinical adoption.
Sub-Sampling
Sub-sampling, a technique for selecting a representative subset of $m$ samples from a larger dataset of size $n$ (where $n\gg m$), has found widespread application in medical image processing (De Boor 1962; Zunair, Rahman, and Mohammed 2019). Specifically, in the context of CT volume analysis for screening tasks, prior research has commonly employed interpolation techniques  [^1] While interpolation methods—which may alter slice content through operations like pixel averaging or summation rather than direct selection—do not strictly conform to the definition of sub-sampling, they are often considered generalized sub-sampling approaches due to their congruent goals of data reduction and efficiency enhancement for downstream tasks. and heuristic slice selection strategies for preprocessing. Interpolation methods include linear interpolation (Needham 1959), spline interpolation (De Boor 1962), and projection interpolation (Liauchuk 2019). While heuristic slice selection methods include subset slice selection (Zunair, Rahman, and Mohammed 2019) and even slice selection (Zunair et al. 2020). However, interpolation methods can alter pixel intensities, risking artifacts and the loss of subtle diagnostic information. Similarly, heuristic slice selection may not comprehensively represent the CT volume, potentially omitting critical diagnostic features. Both limitations can compromise screening accuracy.
More recently, sophisticated machine learning-guided sub-sampling strategies have emerged to address the aforementioned challenges, broadly categorized by their aim for representativeness or diversity. Representative sampling aims to select a subset whose data distribution closely mirrors that of the entire dataset, typically by minimizing certain statistical discrepancies. Prominent examples of such discrepancy measures include the Wasserstein distance (Graf and Luschgy 2007), energy distance (Mak and Joseph 2018), uniform design (Zhang et al. 2023a), maximum mean discrepancy (Chen, Welling, and Smola 2012), and generalized empirical $F$-discrepancy (Zhang et al. 2023b). Conversely, diversity sampling endeavors to select a varied subset encompassing a broad spectrum of informative samples, thereby maximizing the information content within the selected subset (Wu et al. 2024). However, most existing methodologies unilaterally focus on either representativeness or diversity, suffer from inefficiency, and neglect pulmonary anatomical structures, making them incompatible with our task.
Uncertainty Quantification
Uncertainty quantification (UQ) is vital for deploying deep learning models in critical applications like medical diagnosis, enabling them to signal low confidence. Based on the source of the modeled uncertainty, uncertainty can be broadly categorized into two main types: aleatoric uncertainty (AU) (Hüllermeier and Waegeman 2021; Wang et al. 2019) and epistemic uncertainty (EU) (Swiler, Paez, and Mayes 2009; Nannapaneni and Mahadevan 2016; Wang et al. 2025). AU describes the inherent uncertainty within the data itself, while EU captures the uncertainty of the model itself. Prominent UQ approaches include Bayesian Neural Networks (Neal 2012; Ngartera, Issaka, and Nadarajah 2024), which learn distributions over weights but often entail complex and computationally intensive training and inference, for instance, through variational inference (Blundell et al. 2015) or Markov Chain Monte Carlo methods. Monte Carlo (MC) Dropout (Gal and Ghahramani 2016) provides a more practical Bayesian approximation by leveraging dropout at test time, but it requires multiple forward passes, increasing the inference latency by a corresponding multiple. Deep integration (Lakshminarayanan, Pritzel, and Blundell 2017) trains multiple independent models and achieves robust UQ, but at the cost of increasing training, inference, and parameter storage by a corresponding multiple. Other notable methods include evidential deep learning (Sensoy, Kaplan, and Kandemir 2018), which aims for single-pass uncertainty, and conformal prediction (Vovk, Gammerman, and Shafer 2005), providing model-agnostic guarantees but whose utility depends on calibration. While these methods offer valuable uncertainty estimates, their computational burden can be prohibitively high (Abdar et al. 2021), which runs counter to our goal of efficient screening.
Methods
Preliminaries
In this paper, the dataset we used contains $N$ samples (i.e., CT volumes), denoted as $D=\{(\bm{x}_{i},\bm{y}_{i})\}_{i\in[N]}$, where $(\bm{x}_{i},\bm{y}_{i})$ represents the combination of each sample $\bm{x}_{i}$ and its corresponding label $\bm{y}_{i}$, with $N$ being the number of samples. Here, the class label $\bm{y}_{i}$ is in one-hot encoded form. Assuming that each CT volume contains $n$ slices. Since our slice selection method operates within each CT volume, we uniformly denote the slices of each volume as $\{\bm{s}_{i}\}_{i\in[n]}$ to simplify notation, without distinguishing between slices from different volumes. This simplification does not affect subsequent methodological descriptions.
Figure 2: The core insight of CSS. We first select $m$ density peaks for representativeness, and then push them away from the instances selected in the previous iteration via an iterative refinement process with a diversity regularizer. The circle with the black dashed line and the solid black circle denote the selected instances from the previous iteration and the current iteration, respectively.
Slice Sub-Sampling
Within each CT volume, CSS aims to select $m(\ll n)$ slices that could represent the whole CT volume so that the performance of a model trained using the selected slices approaches that of the full CT method. The CSS algorithm operates in three main stages: First, we divide the lung into three parts based on its anatomical structure. Second, we learn semantic representations for each CT slice. Third, we perform an iterative selection process based on representativeness and diversity in each lung part.
Lung Structure Dividing
Considering the anatomical structure of the lungs—where the left lung is divided into upper and lower lobes with a volume ratio of approximately $0.4:0.6$, and the right lung into upper, middle, and lower lobes with a volume ratio of approximately $0.25:0.15:0.6$, we divide the entire CT volume into three distinct parts based on the ratios of $0.25:0.15:0.6$. Following this, the total budget of slice selection $m$ is distributed across these parts in accordance with the same proportion. The subsequent sub-sampling process is performed uniformly across all parts, which ensures the semantic coherence of the sub-sampling results.
Representation Learning
To capture high-level semantic information pertinent to lung pathologies, we leverage the powerful feature extraction capabilities of a pre-trained MedCLIP-ViT model (Wang et al. 2022). The image encoder of MedCLIP-ViT maps $\bm{s}_{i}$ onto a $d$-dimensional hypersphere with $L^{2}$ normalization, denoted as $\bm{h}_{i}=f(\bm{s}_{i})\in\mathbb{R}^{d}$, with $\|\bm{h}_{i}\|_{2}=1$. Crucially, MedCLIP is utilized only for inference to extract these features. This one-time process incurs minimal computational cost. A detailed efficiency analysis is provided in the section of efficiency evaluation.
Cluster-based Sub-Sampling
CSS is guided by two core principles, widely recognized in active learning and unsupervised selection literature (Xie et al. 2023; Wang, Lian, and Yu 2022; Shao et al. 2024): representativeness and diversity. Representativeness ensures that the selected slices capture the primary characteristics of the entire CT volume, while diversity minimizes redundancy among the selected slices. The core insight of CSS is shown in Figure 2.
First, We apply K-Means clustering (Forgy 1965; Lloyd 1982) to partition the $n$ slice features $\{\bm{h}_{i}\}^{n}_{i=1}$ into $m(\leq n)$ distinct clusters $C=\{C_{j}\}_{j=1}^{m}$. Let $\{\bm{z}_{i}^{j}\}_{i=1}^{|C_{j}|}$ be the instance in the cluster $C_{j}$, where $|C_{j}|$ is the number of instances in $C_{j}$. The goal is to minimize the within-cluster sum of squares, which measures the compactness of each cluster:
$\mathop{\min}\limits_{C}\sum\limits_{j=1}^{m}\sum\limits_{i=1}^{|C_{j}|}\|\bm{z}_{i}^{j}-\bm{c}_{j}\|^{2}=\mathop{\min}\limits_{C}\sum\limits_{j=1}^{m}\left|C_{j}\right|\mathop{\mathrm{Var}}(C_{j}),$
where $\bm{c}_{j}$ and $\mathrm{Var}(C_{j})$ are the centroid and variance of cluster $C_{j}$, respectively. This optimization problem is solved iteratively using the Expectation-Maximization (EM) algorithm (McLachlan and Krishnan 2007), starting from randomly initialized centroids. Each cluster $C_{j}$ represents a distinct group of visually or semantically similar instances.
Then, we aim to identify the most representative instance within each cluster $C_{j}$. Intuitively, a representative instance should be ”central” or lie in a high-density region, similar to many other instances within its cluster. We estimate the local density $\rho(\bm{z}_{i}^{j},k)$ around each instance $\bm{z}_{i}^{j}$ using its $k$-th nearest neighbor (k-NN) (Orava 2011):
$\rho(\bm{z}_{i}^{j},k)=\frac{k}{n}\frac{1}{A_{d}\cdot D(\bm{z}_{i}^{j},\bm{z}_{i}^{j}|k)^{d}},$
where $D(\bm{z}_{i}^{j},\bm{z}_{i}^{j}|k)=\|\bm{z}_{i}^{j}-\bm{z}_{i}^{j}|k\|_{2}$ is the Euclidean distance from instance $\bm{z}_{i}^{j}$ to its $k$-th nearest neighbor $\bm{z}_{i}^{j}|k$, $A_{d}=\pi^{d/2}/\Gamma(\frac{d}{2}+1)$ is the volume of a unit $d$-dimensional ball, and $\Gamma(\cdot)$ is the Gamma function. Since our goal is to find the density peak (the most representative instance) within each cluster, and only the relative ranking matters, we can directly select the instance that minimizes the distance $D(\bm{z}_{i}^{j},\bm{z}_{i}^{j}|k)$, which serves as our measure of representativeness. However, the complexity of the naive $k$-NN scales linearly with the number of stored instances, making it infeasible for large-scale datasets. We adopt an approximate $k$-NN search based on Navigable Small World graphs with controllable Hierarchy (HNSW) (Malkov 2018). The time complexity analysis is provided in the Appendix.
Selecting density peaks only might lead to redundancy, especially if cluster boundaries fall within high-density zones, causing selected instances from adjacent clusters to be too similar. To counteract this, we introduce an iterative refinement process with a diversity regularizer, denoted by $\Phi(\bm{z}_{i}^{j},t)$ at iteration $t$. Let $Z^{(t)}=\{\widetilde{\bm{z}}^{(1,t)},\widetilde{\bm{z}}^{(2,t)},...,\widetilde{\bm{z}}^{(m,t)}\}$ be the set of $m$ instances selected at iteration $t$, where each $\widetilde{\bm{z}}^{(j,t)}$ is chosen from cluster $C_{j}$. We set $\Phi(\bm{z}_{i}^{j},0)=0$, i.e., $Z^{(0)}$ is selected using only representativeness without regularization. At iteration $t>0$, for any candidate instance $\bm{z}_{i}^{j}$ in cluster $C_{j}$, its diversity contribution increases the farther it is from the currently selected instances in other clusters ($\widetilde{\bm{z}}^{(l,t-1)}$ for $l\neq j$). We quantify $\Phi(\bm{z}_{i}^{j},t)$, $t>0$, which penalizes proximity to other selected instances:
$\Phi(\bm{z}_{i}^{j},t)=\sum\limits_{l\neq j}\frac{1}{\|\bm{z}_{i}^{j}-\widetilde{\bm{z}}^{(l,t-1)}\|_{2}^{\alpha}},$
where $\alpha$ is a sensitivity hyperparameter controlling the strength of the distance penalty. To stabilize the selection process across iterations, we update this regularizer using an Exponential Moving Average (EMA) (Laine and Aila 2016):
$\overline{\Phi}(\bm{z}_{i}^{j},t)=\beta\cdot\overline{\Phi}(\bm{z}_{i}^{j},t-1)+(1-\beta)\cdot\Phi(\bm{z}_{i}^{j},t),$
where $\beta\in(0,1]$ is the momentum parameter.
At each iteration $t>0$, we select the instance $\bm{z}_{i}^{j}$ within each cluster $C_{j}$ that maximizes a combined score $J(\bm{z}_{i}^{j},t)$, which balances representativeness (high inverse average distance) and diversity (low regularization term):
$J(\bm{z}_{i}^{j},t)=\frac{1}{D(\bm{z}_{i}^{j},\bm{z}_{i}^{j}|k)}-\lambda\cdot\overline{\Phi}(\bm{z}_{i}^{j},t),$
where $\lambda\geq 0$ is a hyperparameter balancing the two objectives. Since calculating $J(\bm{z}_{i}^{j},t)$ for each candidate $\bm{z}_{i}^{j}\in C_{j}$ is no longer feasible for large datasets, we only consider $h$ nearest neighbors of instance $\widetilde{z}^{(j,t-1)}$, denoted by $\{\widetilde{z}^{(j,t-1)}|l\}_{l=1}^{h}$. Then we can derive the selected instance $\widetilde{\bm{z}}^{(j,t)}$ in $C_{j}$ at iteration $t$:
$\widetilde{\bm{z}}^{(j,t)}=\mathop{\arg\max}\limits_{\bm{z}_{i}^{j}\in\{\widetilde{z}^{(j,t-1)}|l\}_{l=1}^{h}}J(\bm{z}_{i}^{j},t).$
The iterative process continues for $T$ iterations, and the final set $Z^{(T)}$ constitutes the selected instances. Now we obtain the dataset after CSS $\widetilde{D}=\{(\widetilde{\bm{x}}_{i},\bm{y}_{i})\}_{i\in[N]}$, where $\widetilde{\bm{x}}_{i}$ represents the sampling result of $\bm{x}_{i}$, containing $m$ representative slices. $\widetilde{D}$ are used to train a downstream model for coarse screening of the pulmonary disease in the next section.
Coarse Screening
Let $F$ be the backbone of the deep model with parameters $\theta$, transforming the sample $\widetilde{\bm{x}}_{i}$ into a feature vector, and $G$ be the linear classifier with the parameter matrix $\bm{W}$. The predictive probability vector of $\widetilde{\bm{x}}_{i}$ is given by $\bm{p}_{i}=\sigma(G(F(\widetilde{\bm{x}}_{i};\theta);\bm{W}))$, where $\sigma$ is the softmax function. The model is trained by minimizing the cross-entropy loss $\mathcal{L}_{CE}$ on $\widetilde{D}$:
$\min_{\theta,\bm{W}}\frac{1}{N}\sum\limits^{N}_{i=1}\mathcal{L}_{CE}(\bm{p}_{i},\bm{y}_{i}).$
The model architecture is shown in the Appendix.
Figure 3: Illustration of HUQ. The circles of different colors represent samples from distinct categories. The solid and dashed lines denote the decision hyperplanes generated by the main classifier and the two auxiliary classifiers, respectively. The gray area indicates the region where high-uncertainty samples are located.
Hybrid Uncertainty Quantification
To quantify aleatoric uncertainty (AU), we modify the model described in the previous section by introducing two auxiliary classifiers, $G_{1}$ with the parameter matrix $\bm{W}_{1}$ and $G_{2}$ with $\bm{W}_{2}$, which operate in parallel with the main classifier $G$. All three classifiers share an identical architecture, ensuring our method can be readily applied to various network structures. The core idea is to maximize the predictive discrepancy between $G_{1}$ and $G_{2}$ while ensuring they both remain accurate classifiers. We define the discrepancy $d_{dis}$ using the $L_{1}$-distance, following the theoretical work on domain adaptation (Ben-David et al. 2010; Cho et al. 2022):
$d_{dis}(\bm{p}_{i}^{1},\bm{p}_{i}^{2})=\frac{1}{K}\left(\|\bm{p}_{i}^{1},\bm{p}_{i}\|_{1}+\|\bm{p}_{i}^{2},\bm{p}_{i}\|_{1}+\|\bm{p}_{i}^{1},\bm{p}_{i}^{2}\|_{1}\right),$ (1)
where $K$ is the number of classes. $\bm{p}_{i}^{1}$ and $\bm{p}_{i}^{2}$ are the softmax outputs for a sample $\widetilde{\bm{x}}_{i}$ from $G_{1}$ and $G_{2}$, respectively. To operationalize this, we formulate the following training objective for the auxiliary classifiers:
$\min\limits_{\bm{W}_{1},\bm{W}_{2}}\frac{1}{N}\sum\limits^{N}_{i=1}\left(\mathcal{L}_{CE}(\bm{p}_{i}^{1},\bm{y}_{i})+\mathcal{L}_{CE}(\bm{p}_{i}^{2},\bm{y}_{i})+\mathcal{L}_{dis}(\bm{p}_{i}^{1},\bm{p}_{i}^{2})\right),$
where $\mathcal{L}_{dis}(\bm{p}_{i}^{1},\bm{p}_{i}^{2})=-d_{dis}(\bm{p}_{i}^{1},\bm{p}_{i}^{2})$. In other words, the first two terms are standard cross-entropy losses that enforce classification accuracy. The third term actively pushes the predictions of $G_{1}$ and $G_{2}$ apart. Importantly, the gradients from this objective are detached from the shared backbone network $F$, which ensures that this adversarial training does not corrupt the primary feature representation. By optimizing this objective, the learned discrepancy $d_{dis}$ becomes a principled measure of AU. Samples lying far from the training manifold or near class boundaries will yield high disagreement, signaling inherent data ambiguity (Hüllermeier and Waegeman 2021; Wang et al. 2019; Cho et al. 2022), as shown in Figure 3.
To capture the model’s parameter uncertainty, i.e., epistemic uncertainty (EU) (Swiler, Paez, and Mayes 2009; Nannapaneni and Mahadevan 2016), we leverage the predictive entropy $\mathcal{H}(\widetilde{\bm{x}}_{i})$ from the main classifier $G$, where $\mathcal{H}(\widetilde{\bm{x}}_{i})=-\sum_{j=1}^{K}\bm{p}_{i,j}\log\bm{p}_{i,j}$, and $\bm{p}_{i,j}$ is the $j$-th value of $\bm{p}_{i}$.
We combine these two distinct but complementary uncertainty sources into a final score $U(\widetilde{\bm{x}}_{i})$:
$U(\widetilde{\bm{x}}_{i})=d_{dis}(\bm{p}_{i}^{1},\bm{p}_{i}^{2})+\mathcal{H}(\widetilde{\bm{x}}_{i}).$
Experiments
Table 1: Comparison with other sub-sampling methods. The best performance is bold, and the second-best performance is underlined, except for the method of full CT volume. Each result shows mean accuracy and standard deviation over $5$-fold cross-validation.
Task $1$ Task $2$ Task $3$
Method Accuracy Recall Accuracy Recall Accuracy Recall
Full CT $99.12_{\pm 0.31}$ $99.15_{\pm 0.35}$ $99.35_{\pm 0.41}$ $99.30_{\pm 0.38}$ $92.47_{\pm 0.60}$ $92.10_{\pm 0.65}$
Interpolation Methods
PI $90.36_{\pm 0.34}$ $89.02_{\pm 0.41}$ $88.92_{\pm 0.40}$ $89.00_{\pm 0.30}$ $84.59_{\pm 0.47}$ $83.50_{\pm 0.52}$
LI $78.52_{\pm 2.54}$ $77.89_{\pm 2.67}$ $73.67_{\pm 2.03}$ $71.70_{\pm 2.30}$ $68.81_{\pm 3.80}$ $67.54_{\pm 4.81}$
SI $98.78_{\pm 0.34}$ $98.20_{\pm 0.37}$ $98.33_{\pm 0.42}$ $98.30_{\pm 0.40}$ $87.24_{\pm 0.40}$ $86.94_{\pm 0.37}$
Heuristic Slice Selection Methods
SSS $96.89_{\pm 0.56}$ $96.28_{\pm 0.62}$ $95.23_{\pm 0.97}$ $95.12_{\pm 0.99}$ $89.34_{\pm 0.90}$ $89.67_{\pm 1.24}$
ESS $96.42_{\pm 0.86}$ $96.85_{\pm 0.78}$ $94.29_{\pm 1.37}$ $94.50_{\pm 1.70}$ $88.56_{\pm 1.56}$ $86.78_{\pm 1.65}$
Representative and Diverse Sub-Sampling Methods
CoreSet $97.38_{\pm 0.47}$ $96.39_{\pm 0.59}$ $98.06_{\pm 0.47}$ $97.57_{\pm 0.77}$ $88.91_{\pm 1.88}$ $89.07_{\pm 2.06}$
ActiveFT $98.75_{\pm 0.38}$ $98.26_{\pm 0.76}$ $98.62_{\pm 0.36}$ $98.38_{\pm 0.61}$ $90.68_{\pm 0.71}$ $90.35_{\pm 1.14}$
CSS (Ours) $\mathbf{98.89_{\pm 0.17}}$ $\mathbf{99.05_{\pm 0.18}}$ $\mathbf{98.65_{\pm 0.23}}$ $\mathbf{98.82_{\pm 0.29}}$ $\mathbf{91.88_{\pm 0.43}}$ $\mathbf{90.85_{\pm 0.54}}$
Dataset and Implementation Details
We conduct comparative experiments on two publicly available datasets: SARS-CoV-2 dataset (Zhang et al. 2020) and LUNG-PET-CT-Dx (Li et al. 2020), which are obtained from the China Consortium of Chest CT Image Investigation (CC-CCII) and The Cancer Imaging Archive (TCIA) (Clark et al. 2013), respectively. The dataset consists of $2,654$ cases in total, including: $311$ cases diagnosed with adenocarcinoma (AC), $747$ cases with Novel Coronavirus Pneumonia (NCP, i.e., COVID-19), $741$ cases with Common Pneumonia (CP) and $855$ normal cases. We organize them into three binary classification tasks. For each task, we evenly divide the dataset into $5$ parts and use $5$-fold cross-validation to evaluate the mean accuracy and its standard deviation. In addition, considering that recall is a crucial indicator in screening tasks, we will also use recall as an evaluation metric. More details of the dataset and the implementations are available in the Appendix.
Comparison with Other Sub-Sampling Methods
In three diagnostic tasks, we compare CSS with different sub-sampling methods, including three widely-used interpolation methods: projection interpolation (PI) (Liauchuk 2019), linear interpolation (LI) (Needham 1959), and spline interpolation (SI) (De Boor 1962), two heuristic slice selection methods: subset slice selection (SSS) (Zunair, Rahman, and Mohammed 2019) and even slice selection (ESS) (Zunair et al. 2020), two representative and diverse sub-sampling methods: CoreSet (Sener and Savarese 2018) and ActiveFT (Xie et al. 2023), and full CT volume method.
For a fair comparison, we set the number of slices after sub-sampling to $64$ for each method, except for the full CT method. Since the full CT method utilizes the complete CT data (differing from other approaches) and serves solely as a baseline reference, we have highlighted it in gray in Table 1. We then use the CT data processed by different sub-sampling methods to train the classification model, while ensuring consistent model architecture and training parameters. Finally, we compare the performance of classification models applied with different sub-sampling methods. We kindly refer the readers to the Appendix for more details about the above methods. The results are shown in Table 1, from which we observe that: (1) The performance of CSS generally surpasses that of other methods, and is very close to that of the full CT volume method across three tasks; (2) Compared to tasks $1$ and $2$, the performance of all methods shows a significant decline in task $3$. This could be attributed to the fact that tasks $1$ and $2$ involve diagnosing pneumonia, where the lesions are larger and more distinct, making them easier to differentiate from normal cases. Task $3$ focuses on detecting adenocarcinoma, where the lesions are smaller and have a more complex morphology, making accurate detection challenging. The visualization results in Figure 4 also prove this point.
Table 2: Comparison with optimal methods of other types under lower sub-sampling rate.
Task Method Accuracy Recall
Task $1$ SI $97.99_{\pm 0.85}$ $98.06_{\pm 0.54}$
SSS $95.38_{\pm 0.46}$ $94.78_{\pm 0.56}$
ActiveFT $98.00_{\pm 0.48}$ $97.32_{\pm 0.47}$
CSS (Ours) $\mathbf{98.13_{\pm 0.66}}$ $\mathbf{98.13_{\pm 0.30}}$
Task $2$ SI $96.73_{\pm 0.95}$ $97.10_{\pm 0.79}$
SSS $94.11_{\pm 0.90}$ $94.06_{\pm 0.89}$
ActiveFT $97.18_{\pm 0.22}$ $97.17_{\pm 0.56}$
CSS (Ours) $\mathbf{97.68_{\pm 0.36}}$ $\mathbf{98.11_{\pm 0.88}}$
Task $3$ SI $83.63_{\pm 1.24}$ $83.49_{\pm 1.33}$
SSS $76.36_{\pm 1.68}$ $77.16_{\pm 1.86}$
ActiveFT $84.73_{\pm 1.15}$ $85.21_{\pm 1.29}$
CSS (Ours) $\mathbf{90.03_{\pm 0.40}}$ $\mathbf{90.35_{\pm 1.14}}$
In addition, we compare CSS with the optimal methods of other types (i.e., SI, SSS, and ActiveFT), and the full CT method under a lower sub-sampling rate setting, that is, selecting $32$ slices from each volume. The experimental results are presented in Table 2, showing that our method achieves over $90\%$ accuracy and recall across all three tasks, outperforming other methods. Especially in complex adenocarcinoma screening tasks (Task $3$), the performance improvement of our proposed CSS compared to the second-best ActiveFT is even more pronounced. CSS outperforms the second-best ActiveFT approach by $5.30\%$ and $5.14\%$ in accuracy and recall, respectively. More comparative experiments at different sub-sampling rates (e.g., sub-sampling to $16$, $8$, or even $4$ slices) are provided in the Appendix.
Table 3: Comparison with other UQ methods.
Method Task $1$ Task $2$ Task $3$
Entropy $50.79_{\pm 9.44}$ $67.74_{\pm 10.78}$ $57.95_{\pm 4.72}$
Discrepancy $74.76_{\pm 15.91}$ $77.18_{\pm 10.25}$ $77.69_{\pm 7.72}$
DE $73.73_{\pm 16.41}$ $69.60_{\pm 8.11}$ $67.56_{\pm 8.89}$
MC Dropout $70.95_{\pm 4.18}$ $78.65_{\pm 5.88}$ $66.28_{\pm 8.07}$
HUQ (Ours) $\mathbf{94.92_{\pm 7.05}}$ $\mathbf{91.59_{\pm 7.93}}$ $\mathbf{92.05_{\pm 5.45}}$
Comparison with Other UQ Methods
We compare our proposed HUQ with four other UQ methods, including two baselines: the entropy-only method and the discrepancy-only method, defined as (1), along with two widely-used UQ methods: Deep Ensemble (DE) (Lakshminarayanan, Pritzel, and Blundell 2017) and Monte Carlo (MC) Dropout (Gal and Ghahramani 2016).
For all test set samples, we calculate uncertainty using different methods and select the top $15$ samples with the highest uncertainty. We then count the number of false negatives ($FN$) and false positives ($FP$) among these samples, denoted as $FN^{\prime}$ and $FP^{\prime}$ respectively. The misdiagnosis recall rate is computed as $Recall_{mis}=(FN^{\prime}+FP^{\prime})/(FN+FP)$. A higher $Recall_{mis}$ indicates that the UQ method more accurately identifies samples with uncertain predictions, and conversely. It is noteworthy that, for a fair comparison, we ensure the classification model’s predictions are consistent before implementing the UQ methods. This is done to guarantee that the denominator part of the $Recall_{mis}$ formula, i.e., $FN+FP$, remains uniform across evaluations. More implementation details are provided in the Appendix. Comparative results are shown in Table 3, from which we draw several observations: (1) Our proposed HUQ achieves the highest $Recall_{mis}$, which confirms that HUQ accurately identifies misdiagnosed cases; (2) Evaluating prediction uncertainty using either AU or EU fails to reliably identify misdiagnosed cases, particularly for entropy-based methods. This is likely because deep models tend to produce overconfident predictions, leading to unreliable final estimations. Thus, the simultaneous consideration of both AU and EU estimation is essential.
Efficiency Evaluation
We analyze the efficiency of the sub-sampling methods and the UQ methods, respectively. For a more visual comparison, we demonstrate the actual runtime of them here. The time complexity analysis is shown in the Appendix.
Sub-Sampling
We calculate the inference times of CSS, the optimal methods of other types (i.e., SI, SSS, and ActiveFT), and the full CT method. The inference time of ActiveFT and CSS comprises three components: (1) Feature extraction time using the pre-trained MedCLIP-ViT; (2) Time taken to select $32$ slices based on their respective strategies; (3) Time for disease detection by the classification model. The inference time of the SI and SSS methods only includes the time for slice selection (or interpolation) and model classification. The full CT method’s computation time only includes the inference time of the classification model. For fair comparison, all methods are evaluated under identical hardware configurations, and none include the UQ procedure. As shown in Table 4, although CSS has a longer inference time compared to SI and SSS methods, these two methods exhibit significant performance degradation at lower sampling rates, especially in Task $3$, thus cannot guarantee accurate screening. Compared to the full CT method, CSS reduces the inference time by over $60\%$.
Table 4: Inference time for $50$ cases, with each containing $32$ slices. The percentage in brackets represents the proportion of time consumed relative to that of the full CT method.
Method Task $1$ Task $2$ Task $3$
SI $52.83s\thinspace(31.22\%)$ $55.35s\thinspace(29.28\%)$ $53.68s\thinspace(29.95\%)$
SSS $41.52s\thinspace(24.53\%)$ $42.78s\thinspace(22.63\%)$ $41.76s\thinspace(23.30\%)$
ActiveFT $81.28s\thinspace(48.03\%)$ $87.57s\thinspace(46.32\%)$ $82.37s\thinspace(45.95\%)$
CSS (Ours) $66.64s\thinspace(39.38\%)$ $70.88s\thinspace(37.49\%)$ $67.53s\thinspace(37.67\%)$
Full CT $169.24s$ $189.05s$ $179.25s$
Table 5: Extra time consumed by different UQ methods relative to the multiple of one inference time. $M_{DE}$ represents the number of models that DE ensembles, and $M_{MC}$ represents the number of MC Dropout forward passes.
Method Extra time consumption
Entropy 0$\approx$ 0$\times$ 0 ×
Discrepancy 0$\approx$ 0$\times$ 0 ×
DE $\approx(M_{DE}-1)\times$
MC Dropout $\approx(M_{MC}-1)\times$
HUQ (Ours) 0$\approx$ 0$\times$ 0 ×
Uncertainty Quantification
We calculate the extra latency of different UQ methods as multiples of single-inference time. Results in Table 5 reveal that compared to model inference time, our proposed HUQ adds negligible computational overhead. This efficiency stems from its shared backbone outputs and the fact that forward propagation through the additional classification heads incurs minimal latency relative to the backbone. Similarly, both entropy-only and discrepancy-only methods introduce virtually no extra computation time. DE, however, requires multiple model predictions and thus multiplies the inference latency accordingly. MC Dropout needs repeated forward passes, proportionally increasing latency. Consequently, our approach achieves simultaneous estimation of AU and EU while maintaining near-zero computational overhead.
Figure 4: Visualization of CT slices selected by our proposed CSS. The red, yellow, and blue boxes mark the lesion areas of NCP, CP, and AC, respectively.
Visualization
We visualize the CT slices selected by CSS across three tasks, as shown in Figure 4, where each task displays one case, with each case retaining $4$ slices. From the visualization results in Figure 4, we have several observations: (1) the selected slices contain lesion areas, indicating that the CSS method can select representative slices to aid in diagnosis; (2) Compared to AC, the lesion areas of NCP and CP are larger and occur more frequently, which indicates that AC is less obvious on CT and more difficult to diagnose, consistent with the results in Tables 1 and 2. More visualization results are provided in the Appendix.
Conclusions
In this work, we present an efficient and reliable framework for pulmonary disease screening to address the dual challenges of computational cost and model trustworthiness in analyzing large-scale CT volumes. Our proposed CSS method efficiently curates a small yet highly informative subset of slices. To enhance clinical safety and interpretability, we introduced a novel HUQ mechanism that efficiently estimates both aleatoric and epistemic uncertainty with minimal overhead. Our extensive experiments on two public datasets demonstrate that the proposed framework achieves diagnostic performance comparable to computationally intensive full-volume analysis while reducing inference time by over $60\%$. Crucially, our method’s ability to reliably flag uncertain cases for expert review marks a significant step toward the deployment of safe, efficient, and trustworthy AI in real-world clinical workflows.
Acknowledgment
We would like to thank Yiqun Fan for the timely assistance provided in analyzing the selected CT slices during the visualization experiments.
Appendix A Model Architecture of Coarse Screening
First, we employ a ResNet-34 (He et al. 2016) as the backbone for extracting features from each individual selected slice. We modify the standard ResNet-34 by replacing the final $1000$-dimensional fully-connected (FC) layer with a $512$-dimensional FC layer and removing the final softmax function. Then, the output of ResNet-34 is used as input to a light transformer-based network (Vaswani et al. 2017) for feature fusion, where both the layer number and the head number of the Transformer Encoder are set to $6$. Finally, we perform an FC layer and a softmax function on the output of the network to predict the probability of disease.
To ensure fair comparison, we configure the same classification model for all sub-sampling methods and compare the performance of models trained with data sub-sampled by different methods. Although the feature extraction process is at the slice level, we fuse them, and the model prediction output is at the patient level, so there is no data leakage problem, that is, slices from the same patient could not appear in both the training set and the validation set. In addition, it is worth noting that the full CT method uses 3D-ResNet (Fan et al. 2021) as the classification model, therefore, there will be no data leakage problem.
Appendix B Preprocessing and Division Details of Datasets
Given our objective to sub-sample CT slices, we eliminate cases with a relatively low number of slices, specifically those with less than $64$ slices. After that, $2,654$ cases are included in the study, containing $311$ cases diagnosed with adenocarcinoma (AC), $747$ cases with Novel Coronavirus Pneumonia (NCP, i.e., COVID-19), $741$ cases with Common Pneumonia (CP) and $855$ normal cases. Here, a case refers to a 3D CT volume. The AC cases are from LUNG-PET-CT-Dx, while the NCP, CP and normal cases are from the SARS-CoV-2. Then, we design three binary classification tasks to detect the above three diseases. Specifically, Task $1$ contains $747$ NCP and $855$ normal cases; Task $2$ contains $741$ CP and $855$ normal cases; Task $3$ contains $311$ AC and $311$ normal cases. The normal cases for Task $3$ are randomly chosen from $855$ normal cases of the SARS-CoV-2.
Appendix C Implementations Details
Computer Resources
All models are implemented in PyTorch and trained on an RTX 4090 with $24$ GB memory.
Sub-Sampling
CSS: The CSS process is described in the main text, so we will not reiterate it here. Below, we provide the specific hyperparameter settings as follows: $k=10$, $\alpha=0.5$, $\beta=0.9$, $\lambda=0.5$, and $T=10$. It is worth noting that in the CSS method, the distance between each instance and its $k$-th nearest neighbor can be sensitive to noise, as it relies solely on the $k$-th nearest neighbor. Following the settings of (Wang, Lian, and Yu 2022), we modify this by using the average distance to the $k$ nearest neighbors, denoted as $\overline{D}(\bm{z}_{i}^{j},k)=\frac{1}{k}\sum_{l=1}^{k}D(\bm{z}_{i}^{j},\bm{z}_{i}^{j}|l)$.
CoreSet (Sener and Savarese 2018) and ActiveFT (Xie et al. 2023): For fair comparison, we set the preprocessing steps (i.e., feature extraction) and the lung patching process of the CoreSet and ActiveFT to be consistent with the CSS, and then perform slice selection according to their respective strategies. Their respective strategies can be found in the original paper, and are not described here again.
PI, LI and SI: Interpolation methods calculate the new pixel value in the compressed image based on the values of surrounding pixels on the $Z$-axis.
SSS: SSS selects $m/3$ slices from the first, middle, and last positions of the CT volume, respectively.
ESS: ESS selects one after every $\lfloor n/m\rfloor$ of slices.
Classification Model Training
During the training stage of the classification model, we use adaptive moment estimation (Adam) with an initial learning rate $1e^{-3}$ to optimize the network. We set different batch sizes for experiments with different numbers of selected slices. The more slices selected for each case, the more GPU memory each case occupies; thus, the batch size should be smaller. In detail, $batchsize=1,4,8,16,16$ is set for $m=64,32,16,8,4$ respectively. For a fair comparison, the training epoch of all experiments is set to $400$.
Uncertainty Quantification
To ensure a fair comparison of the performance of UQ methods, the classification prediction outcomes of the models must remain consistent. Some UQ methods produce multiple outputs, such as the DE methods, which integrate predictions from multiple models, and the discrepancy-only methods, which employ multiple classifiers. To address this, we establish a primary model (i.e., ResNet-34) for all methods, ensuring identical initialization and training configurations. Specifically, this model is trained using CT data sub-sampled via CSS, with the number of sub-sampled slices set to $32$. This approach guarantees consistent classification predictions across all methods while allowing different approaches to incorporate outputs from auxiliary branches during the UQ phase.
Entropy-only method: The entropy of the model’s predicted probability vector is used as the uncertainty measure. For sample $\widetilde{\bm{x}}_{i}$, entropy is calculated as: $\mathcal{H}(\widetilde{\bm{x}}_{i})=-\sum_{j=1}^{K}\bm{p}_{i,j}\log\bm{p}_{i,j}$, where $\bm{p}_{i,j}$ denotes the $j$-th element of $\bm{p}_{i}$.
Discrepancy-only method: This method employs the same model structure and training protocol as HUQ, but utilizes only discrepancy as the uncertainty metric. The discrepancy calculation is denoted as (1).
DE method: Two auxiliary deep models are trained. Their architectures and training settings match the primary model, but with different initializations. During the UQ phase, the predictive probabilities of all three models are averaged, after which entropy is computed as the uncertainty measure.
MC Dropout method: An additional neural network with Dropout layers is trained. It shares the same architecture as the primary model but includes Dropout layers (rate=$0.5$) after each block. During the UQ phase, this model performs two additional forward passes. Its probabilistic outputs are averaged with those of the primary model, and entropy is subsequently computed as the uncertainty measure.
Appendix D Time Complexity
Sub-Sampling
We analyze the asymptotic time complexity of our proposed CSS, the optimal sub-sampling methods of other types, and the full CT method. The results are shown in Table 6, from which we observe that: (1) The time complexity of CSS is higher than that of the full CT method. However, the full CT method is based on 3D networks, and 3D convolutions involve far more floating-point operations than 2D convolutions. Thus, the actual runtime of the full CT method significantly exceeds that of CSS. (2) Compared with ActiveFT, which achieves suboptimal performance across different sampling rates, CSS also exhibits lower time complexity. This is because we employ HNSW to improve efficiency when computing pairwise sample similarities (the dominant computational bottleneck).
Table 6: Time complexity of different sub-sampling methods. $N$ and $n$ are the number of CT volumes and the number of slices per volume, respectively.
Method Time complexity
Full CT $O(N\cdot n)$
SI $O(N\cdot n)$
SSS $O(N\cdot m)$
ActiveFT $O(N\cdot n^{2})$
CSS (Ours) $O(N\cdot n\cdot\log n)$
Uncertainty Quantification
We analyze the asymptotic time complexity of our proposed HUQ and four other UQ methods. The results are presented in Table 7, which shows that compared to the DE and MC Dropout methods, our proposed HUQ, entropy-only, and discrepancy-only methods exhibit lower time complexity, while HUQ additionally maintains both AU and EU.
Table 7: Time complexity of UQ methods. $M_{DE}$ represents the number of models that DE ensembles, and $M_{MC}$ represents the number of MC Dropout forward passes.
Method Time complexity
Entropy $O(N\cdot n)$
Discrepancy $O(N\cdot n)$
DE $O(N\cdot n\cdot M_{DE})$
MC Dropout $O(N\cdot n\cdot M_{MC})$
HUQ (Ours) $O(N\cdot n)$
Table 8: Comparison under extreme sub-sampling.
Task $1$ Task $2$ Task $3$
Number Method Accuracy Recall Accuracy Recall Accuracy Recall
$16$ SI $96.99_{\pm 1.55}$ $97.21_{\pm 1.10}$ $93.87_{\pm 2.27}$ $94.07_{\pm 2.03}$ $76.71_{\pm 2.70}$ $76.90_{\pm 2.56}$
ActiveFT $96.94_{\pm 0.87}$ $97.19_{\pm 0.88}$ $93.61_{\pm 1.48}$ $94.20_{\pm 1.95}$ $78.95_{\pm 2.11}$ $79.43_{\pm 2.03}$
CSS (Ours) $\mathbf{97.26_{\pm 1.54}}$ $\mathbf{97.58_{\pm 1.21}}$ $\mathbf{94.00_{\pm 2.15}}$ $\mathbf{94.28_{\pm 2.04}}$ $\mathbf{85.60_{\pm 2.37}}$ $\mathbf{85.54_{\pm 2.40}}$
$8$ SI $95.96_{\pm 1.19}$ $95.88_{\pm 1.22}$ $92.09_{\pm 2.77}$ $92.55_{\pm 2.49}$ $67.34_{\pm 3.83}$ $67.68_{\pm 3.41}$
ActiveFT $96.13_{\pm 1.11}$ $96.12_{\pm 1.43}$ $92.42_{\pm 1.75}$ $92.98_{\pm 2.28}$ $74.61_{\pm 3.13}$ $75.24_{\pm 2.41}$
CSS (Ours) $\mathbf{96.57_{\pm 1.24}}$ $\mathbf{96.89_{\pm 1.27}}$ $\mathbf{93.21_{\pm 3.72}}$ $\mathbf{93.02_{\pm 3.44}}$ $\mathbf{82.55_{\pm 2.32}}$ $\mathbf{82.63_{\pm 2.17}}$
$4$ SI $93.86_{\pm 2.24}$ $94.24_{\pm 1.89}$ $90.03_{\pm 2.25}$ $90.78_{\pm 2.30}$ $64.41_{\pm 3.70}$ $64.57_{\pm 3.54}$
ActiveFT $94.07_{\pm 1.40}$ $94.51_{\pm 2.02}$ $90.79_{\pm 2.08}$ $91.23_{\pm 2.44}$ $70.72_{\pm 3.41}$ $69.14_{\pm 2.98}$
CSS (Ours) $\mathbf{94.90_{\pm 2.12}}$ $\mathbf{95.19_{\pm 2.01}}$ $\mathbf{91.17_{\pm 2.54}}$ $\mathbf{91.76_{\pm 2.51}}$ $\mathbf{78.64_{\pm 3.19}}$ $\mathbf{78.89_{\pm 2.98}}$
Figure 5: Additional visualizations of CT slices selected by our proposed CSS. The red, yellow, and blue boxes mark the lesion areas of NCP, CP, and AC, respectively.
Appendix E Additional Comparison with Other Sub-Sampling Methods
We compare CSS with SI and ActiveFT under extreme sub-sampling, that is, selecting $16$, $8$, and $4$ CT slices from each case. The experimental results are shown in Table 8, from which we can observe that our method outperforms other methods, especially on the difficult task $3$. For instance, at $Number=4$, CSS achieved an accuracy of $78.64\%$ and a recall of $78.89\%$. Although this shows a significant drop compared to results with higher sampling rates, it surpasses the second-best method ActiveFT by $7.92\%$ in accuracy and $9.75\%$ in recall. This demonstrates that under extreme sub-sampling scenarios, our method holds a greater advantage.
Appendix F Additional Visualization Results
We show additional visualization results as shown in Figure 5. Each task includes $2$ cases, and each case contains $4$ slices selected by our proposed CSS. We can observe that AC lesions appear less frequently in the slices compared to NCP and CP, which means it is more difficult to be accurately screened, especially in the early stages of the disease. Our CSS method is able to select slices containing lesion information even under a very low sampling rate, i.e., sub-sampling to $4$ slices.